A bag contains equal number of triangles,squares,rectangles,pentagons and trapezia. If one shape is picked from the bag what is the probability that the shape is 
1.A square
2.four sided
3.not a triangle 
 
 

Grade:10

1 Answers

Sandeep Bhamoo
5 years ago
Probability(square) =1/5, probability(4 sided)=3/5. Probability (not a trngl)=4/5. Please do approve my answer if correct